A 25.5 kg person climbs up a uniform ladder
with negligible mass. The upper end of the
ladder rests on a frictionless wall. The bottom
of the ladder rests on a floor with a rough
surface where the coefficient of static friction
is 0.22 . The angle between the horizontal and
the ladder is θ . The person wants to climb
up the ladder a distance of 0.98 m along the
ladder from the ladder’s foot.What is the minimum angle θmin (between
the horizontal and the ladder) so that the
person can reach a distance of 0.98 m without
having the ladder slip? The acceleration of
gravity is 9.8 m/s
2.Answer in units of ◦. Don't round answer.
1
Expert's answer
2020-01-21T03:32:57-0500
"W=(25.5)(9.8)=249.9\ N"

"F_f=(0.22)(249.9)=54.978\ N"

Counter clockwise torque:


"249.9(0.98)\sin\theta"

Clockwise torque:

"54.978(0.98)\cos\theta"

They need to be equal.


"\tan{\theta}=\frac{54.978}{249.9}"

"\theta=12.4\degree"